# On the Effects of Data Scale on Computer Control Agents

### v3

@@ Abstract
Autonomous agents that control computer interfaces to accomplish human tasks are emerging. Leveraging LLMs to power such agents has been of special interest, but unless fine-tuned on human-collected task demonstrations, performance is still relatively low. In this work we study whether fine-tuning alone is a viable approach for building real-world computer control agents. To this end we collect and release a new dataset, AndroidControl, consisting of 15,283 demonstrations of everyday tasks with Android apps. Compared to existing datasets, each AndroidControl task instance includes both high and low-level human-generated instructions, allowing us to explore the level of task complexity an agent can handle. Moreover, AndroidControl is the most diverse computer control dataset to date, including 15,283 unique tasks over 833 Android apps, thus allowing us to conduct in-depth analysis of the model performance in and out of the domain of the training data. Using the dataset, we find that when tested in domain fine-tuned models outperform zero and few-shot baselines and scale in such a way that robust performance might feasibly be obtained simply by collecting more data. Out of domain, performance scales significantly more slowly and suggests that in particular for high-level tasks, fine-tuning on more data alone may be insufficient for achieving robust out-of-domain performance.
@@ 1 Introduction
Recent work has studied how large language models (LLMs) can be leveraged to build computer control agents [17, 42, 38, 16, 6] that accomplish human tasks by interacting with a computer environment. These agents perceive the state of the computer by observing its screen (from screenshots or application UI trees), and generate actions (click, type, scroll, etc.) that are executed through the computer’s user interface. Tasks, specified in natural language, can range from configuring device settings and sending emails, to navigating shopping websites and planning a trip.
While progress is rapidly advancing, absolute performance of computer control agents that leverage pre-trained LLMs without fine-tuning on task demonstrations is still relatively low. When tested in real-world environments, where agents control everyday applications and websites, recently-reported task success rates range from 12% on desktop applications [37] to 46% on mobile applications [3]. In contrast, agents that leverage models fine-tuned for task execution [24, 9], achieve even 80% [9] success rate, when tested on websites and tasks similar to what they are trained on.
While the pattern of collecting new datasets and fine-tuning shows promise, there are at least two important unanswered questions. First, to the best of our knowledge no prior work has examined the question of scaling: how much data must be collected in order to obtain a given performance level with fine-tuned models. This question is particularly important because human demonstrations of computer interactions for fine-tuning are time consuming and expensive to collect. Understanding how performance scales, both in domain and out of the domain of the collected demonstrations (unseen tasks and unseen applications), is important for determining whether fine-tuning alone is a viable path towards deploying computer control agents in the real world. Therefore, one of the main goals of this work is to rigorously quantify how performance of fine-tuned agents scales, both in and out of domain, as the amount of data used for fine-tuning is increased.
Second, it is not clear the level of task complexity fine-tuning might be fruitfully applied to. Conceptually, computer control agents must both decompose a high-level goal into a set of small atomic actions and execute (“ground”) those actions in a device screen. While the high-level reasoning with LLMs, required for determining how to accomplish high-level goals, is still an open problem in artificial intelligence [34, 33, 41, 44, 18, 40, 35], the set of low-level actions (clicking, typing, etc…) required to execute tasks are more constrained, and general agents capable of robust grounding across domains might be approachable via fine-tuning. Therefore, a second goal of this work is to quantify the scaling of fine-tuning for agents performing both high-level and low-level tasks.
@@
Rigorously quantifying scaling in these ways requires a carefully constructed dataset. To this end, we introduce AndroidControl, a large-scale dataset of 15,283 demonstrations of tasks performed by humans in Android apps. Figure 1 shows an example data sample. Compared to existing datasets [6, 26], for every task AndroidControl provides both the high- and low-level human-generated instructions describing it. This is essential to investigate the level of task complexity a model can handle and also provides richer supervision during training. AndroidControl is also the most diverse UI control dataset that exists today, including 15,283 unique tasks over 833 different Android apps, thus allowing us to generate multiple test splits for measuring performance in and out of domain. As a resource to the community, we make AndroidControl publicly available. [^1] https://github.com/google-research/google-research/tree/master/android_control
Overall, we make the following contributions: (i) we collect and release AndroidControl, a new computer control dataset whose size, structure and diversity advances previous datasets, (ii) we use AndroidControl to quantify how fine-tuning with demonstrations scales when applied to both low- and high-level tasks and to tasks in and out of the domain of the training data, and (iii) we compare fine-tuning to various zero-shot and few-shot baselines, finding that fine-tuning scales favorably in domain, but out of domain, it requires one or two orders of magnitude more data to obtain robust performance on high-level tasks, suggesting that additional approaches may be beneficial for obtaining agents which robustly perform out-of-domain high-level tasks.
@@ 2 Related work
Table 1: Comparison of AndroidControl to existing datasets. We consider device platform, size (as number of task demonstrations), diversity (as number of unique task instructions and apps/websites), average number of steps in a task, how the UI state is captured (UI tree vs. screenshot), and whether tasks are described through high-level instructions or sequences of low-level commands.
@@
Table 1 compares AndroidControl to existing computer control datasets. The structure of these datasets is similar. They consist of a natural language task description and a human-recorded demonstration, in the form of a sequence of UI actions (click, type, swipe, etc.) and associated UI states. What differentiates these datasets is mainly whether they are single-step (as in grounding referring expressions datasets like UIBert [2]), whether the task description is expressed as a high-level goal or as a sequence of low-level step instructions, and how the UI state is represented (screenshot vs. UI tree). Three features make AndroidControl unique. First, for every task, it contains both low-level and high-level instructions generated by human annotators. While a few other datasets contain both these types of annotation, their low-level instructions are either synthetically generated (as in MoTIF [5]) or are limited to one action type (only click actions [21, 32]). In addition to bringing richer language supervision during training, the availability of human-generated low-level instructions allows us to test computer control agents on different levels of task complexity. Second, if we consider the number of unique task instructions and the number of human demonstrations, AndroidControl is the second-largest dataset to date, second only to AitW [26]. However, AitW does not contain application UI trees, thus making the UI state representations incomplete, and despite its size it covers a much smaller set of apps. The diversity of task scenarios present in AndroidControl is in fact its third differentiating feature: AndroidControl includes tasks from 833 different Android apps, 6 times more than popular datasets like Mind2Web [6] and 2 times more than AitW. This diversity makes AndroidControl optimal for realistic, out-of-domain analysis. Note that Mind2Web also provides out-of-domain splits but given its smaller size (2,350 tasks over 137 websites, with a train split of 1k demonstrations) is not suitable for a scaling analysis.
@@
In addition to the datasets listed in Table 1, recent work proposes interactive testing environments for computer control agents [30, 45, 17, 37, 27] where the environment provides the agents with reward signals. These environments are designed for online testing and are limited to no more than 20 applications or websites. The only exception is MiniWob [30] for which task demonstrations have been collected, but the environment consists of much simplified, synthetic websites.
Computer control agents
@@
Early computer control agents were trained from scratch using behavioural cloning [14] or reinforcement learning [22, 10]. Current computer agents use pre-trained LLMs and multimodal models. One line of work prompts LLMs in a zero-shot or few-shot regime [38, 26, 11, 17, 16, 43]. Another line of work relies on fine-tuning which is applied end to end [24] or to build specific model capabilities, such as identifying the interactable UI elements in a webpage [42, 9, 6]. To name a few, SeeAct [42], which we use in our evaluation, is a web agent that leverages large multimodal models to understand text and visual elements on webpages. The best-performing SeeAct agent relies on a fine-tuned cross-encoder model to select candidates web elements for interaction. WebGPT [24] fine-tunes GPT-3 to learn to use a web browser. WebAgent [9] pre-trains a T5 model to extract HTML snippets and leverages Flan-U-PaLM to generate Python code to control a web environment. Synapse [43] introduces a trajectory-as-exemplar prompting method where memory of previous interactions allows the agent to perform complex, multi-step tasks.
@@
As evidenced by various LLM studies [4, 15, 12], scaling model and data size for the training leads to steady improvements in domain generalization. On the other hand, when transferring a pre-trained model to a downstream task through fine-tuning, while in-distribution performance improves, a reduction in the robustness to distribution shifts is observed [19, 36, 1]. In this work, we empirically study how scaling data size in fine-tuning affects in-domain and out-of-domain performance of computer control agents. While prior work has tested computer agents using out-of-domain test splits [6, 5], to the best of our knowledge a data scale analysis has not been conducted. To minimize training cost and to maintain the out-of-domain generalization, we evaluate also the option of not fine-tuning, by evaluating multiple zero-shot and few-shot baselines.
@@ 3.1 Data collection
During collection of a demonstration, annotators first provide a high-level description of a task in natural language (e.g., "Add an alarm to wake me up on Saturday mornings at 6am"). We ask annotators to make the descriptions detailed enough to be interpretable without any ambiguity. We also instruct them to always include the name of the target app in the task description, unless obvious (e.g., Google first-party apps such Clock or Settings). By doing so, the collected data can enable us to test memory-less, single-turn agent interactions.
@@ 3.3 Dataset splits
We create a train, a validation and 4 test splits whose number of task demonstrations (episodes) and characteristics are detailed in Table 3. In order to measure how performance scales in domain and out of the domain of the collected data, we create the following test sub-splits: 1) in domain data (IDD): randomly pulled episodes from the same distribution as the training data; 2) unseen-app: a test split using apps not present in the train split; 3) unseen-task: a test split with tasks not present in the train split; and 4) unseen-category: a test split with apps from categories not present in the train split. Note that the test splits may contain overlapping episodes. For example, episodes in the unseen-category split will also be in the unseen-app and unseen-tasks splits.
@@ 4.1 Agent implementation
We implement a computer control agent for Android. The agent receives task instructions expressed in natural language. It observes the environment (the device) by deriving textual representations of the screen directly from the Android accessibility tree. The screen representation lists the on-screen UI elements. Each element is described according to the following attributes: type, text, content description, bounding boxes and various state tags (e.g., clickable, scrollable, focused, checked). As mobile UI screens may contain hundreds of UI elements (200 on average in AndroidControl, Table 3), we pre-process the screen representations to include only UI elements that have a non-empty text description or UI elements of critical types (switch and edit). This process facilitates the agent’s task and reduces the input’s size. Note that our agent implementation does not directly leverage the page screenshot. While recent work explores how to infer screen representations from raw screens [17], best performance is still reported when using accessibility trees or HTML [37, 42, 27]. We expect the general trends we observe will hold true for multimodal language models.
@@
The agent predicts an action among a set of candidate actions. The set of available actions matches the actions defined by AndroidControl (Table 2) with two main changes. We add a terminate action which the agent predicts when it deems the task complete or infeasible. As this action is not originally provided in the dataset, for training purposes, we artificially insert it at the end of every episode (see Appendix D.1). For efficiency reasons, as in prior work [20], the input_text action is modified to include also the preceding click action necessary to set focus on the target element. The agent predicts the action type and any required arguments for the action, specifically the target UI element for a click action, the text to be typed and the target element for a typing action, the name of an app to open, the direction of a scroll, etc. For an example of screen representation, a summary of the agent’s action space, and more details on the agent implementation please refer to Appendix C.
@@ 4.2 Experimental setup
The LLMs we experiment with include PaLM-2L [8], PaLM-2S [8], Gemini 1.5 Pro [7], GPT-4 and GPT-4 Turbo [25]. We set the temperature to zero for all models to obtain more deterministic responses. To limit compute, we perform fine-tuning only with PaLM-2S, and adopt the parameter efficient tuning approach of LoRA [13]. We set the LoRA rank to 4 when fine-tuning with small amounts of data ($<10$k episodes), while switch to a rank of 64 when using more episodes. For few-shot experiments we use Gemini 1.5 Pro which provides a context window of 1M tokens.
We create SeqIO [28] tasks to extract data points from AndroidControl and to generate prompts and target outputs (more details in Appendix D.1). We setup two SeqIO tasks: (i) SeqIO HL (high-level) where only a high-level instruction is included in the prompt, and (ii) SeqIO LL (low-level) where both a low-level instruction and its corresponding high-level instruction are included. This second task emulates the use case where an LLM is used for decomposing high-level instructions into a sequence of low-level commands and another LLM, which is used for grounding, may improve performance by having access to the context of the high-level command. In addition to the NL instruction(s), each data point contains the textual description of the start screen, the history of performed actions, and the ground-truth action. Through these two SeqIO tasks, we investigate how a model performs on simpler (LL) or harder (HL) task instructions.
@@
We test four zero-shot methods. (i) We use the AitW [26] prompt, specifically designed for Android and the PaLM model, without any modifications. (ii) We adapt the best-performing SeeAct [42] variant ("choice") which grounds actions via textual choices. SeeAct was originally designed for GPT-4V for web navigation tasks. At each step, SeeAct queries the LLM twice. In the first query, it analyzes the state and performs reasoning. Then, in the second query, it asks the LLM to select an action from multiple choices. We use the SeeAct prompt by Rawles et a.. [27] adapted to work on mobile and to take textual representations of Android screens as input. (iii) We evaluate the text-only version of M3A [27] that combines ReAct-style [41] and Reflexion-style [31] prompting. (iv) Finally, we test a zero-shot prompt (implementation in Appendix D.2.2) of the same form we use with our agent described above. This allows us to measure performance of a base model for our agent without any fine-tuning. This prompt emphasizes the use of a screen description composed of UI elements, hence the name ER (Element Representations). Note that with the exception of ER, which we ran with all 4 base models, to limit prompt sensitivity [29], we ran the other prompts with the model family they were originally designed for.
@@
Table 4: Performance on the IID sub-split of Random-500. For few-shot (FS) and LoRA-tuned (LT) methods -$X$ (-$all$) indicates $X$ ($all$) episodes are used in the prompt or in training. Unless noted with a “r64” for models fine-tuned with a rank of 64, LoRA-tuned models use rank=4.

### v5

# On the Effects of Data Scale on UI Control Agents

@@ Abstract
Autonomous agents that control user interfaces to accomplish human tasks are emerging. Leveraging LLMs to power such agents has been of special interest, but unless fine-tuned on human-collected task demonstrations, performance is still relatively low. In this work we study whether fine-tuning alone is a viable approach for building real-world UI control agents. To this end we collect and release a new dataset, AndroidControl, consisting of 15,283 demonstrations of everyday tasks with Android apps. Compared to existing datasets, each AndroidControl task instance includes both high and low-level human-generated instructions, allowing us to explore the level of task complexity an agent can handle. Moreover, AndroidControl is the most diverse UI control dataset to date, including 14,548 unique tasks over 833 Android apps, thus allowing us to conduct in-depth analysis of the model performance in and out of the domain of the training data. Using the dataset, we find that when tested in domain fine-tuned models outperform zero and few-shot baselines and scale in such a way that robust performance might feasibly be obtained simply by collecting more data. Out of domain, performance scales significantly more slowly and suggests that in particular for high-level tasks, fine-tuning on more data alone may be insufficient for achieving robust out-of-domain performance.
@@ 1 Introduction
Recent work has studied how large language models (LLMs) can be leveraged to build UI control agents [18, 43, 39, 17, 7] that accomplish human tasks by interacting with a digital device environment. These agents perceive the state of the device by observing its screen (from screenshots or application UI trees), and generate actions (click, type, scroll, etc.) that are executed through the device’s user interface. Tasks, specified in natural language, can range from configuring device settings and sending emails, to navigating shopping websites and planning a trip.
While progress is rapidly advancing, absolute performance of UI control agents that leverage pre-trained LLMs without fine-tuning on task demonstrations is still relatively low. When tested in real-world environments, where agents control everyday applications and websites, recently-reported task success rates range from 12% on desktop applications [38] to 46% on mobile applications [3]. In contrast, agents that leverage models fine-tuned for task execution [25, 10], achieve even 80% [10] success rate, when tested on websites and tasks similar to what they are trained on.
While the pattern of collecting new datasets and fine-tuning shows promise, there are at least two important unanswered questions. First, to the best of our knowledge no prior work has examined the question of scaling: how much data must be collected in order to obtain a given performance level with fine-tuned models. This question is particularly important because human demonstrations of UI interactions for fine-tuning are time consuming and expensive to collect. Understanding how performance scales, both in domain and out of the domain of the collected demonstrations (unseen tasks and unseen applications), is important for determining whether fine-tuning alone is a viable path towards deploying UI control agents in the real world. Therefore, one of the main goals of this work is to rigorously quantify how performance of fine-tuned agents scales, both in and out of domain, as the amount of data used for fine-tuning is increased.
Second, it is not clear the level of task complexity fine-tuning might be fruitfully applied to. Conceptually, UI control agents must both decompose a high-level goal into a set of small atomic actions and execute (“ground”) those actions in a device screen. While the high-level reasoning with LLMs, required for determining how to accomplish high-level goals, is still an open problem in artificial intelligence [35, 34, 42, 45, 19, 41, 36], the set of low-level actions (clicking, typing, etc…) required to execute tasks are more constrained, and general agents capable of robust grounding across domains might be approachable via fine-tuning. Therefore, a second goal of this work is to quantify the scaling of fine-tuning for agents performing both high-level and low-level tasks.
@@
Rigorously quantifying scaling in these ways requires a carefully constructed dataset. To this end, we introduce AndroidControl, a large-scale dataset of 15,283 demonstrations of tasks performed by humans in Android apps. Figure 1 shows an example data sample. Compared to existing datasets [7, 27], for every task AndroidControl provides both the high- and low-level human-generated instructions describing it. This is essential to investigate the level of task complexity a model can handle and also provides richer supervision during training. AndroidControl is also the most diverse UI control dataset that exists today, including 14,548 unique tasks over 833 different Android apps, thus allowing us to generate multiple test splits for measuring performance in and out of domain. As a resource to the community, we make AndroidControl publicly available. [^1] https://github.com/google-research/google-research/tree/master/android_control
Overall, we make the following contributions: (i) we collect and release AndroidControl, a new UI control dataset whose size, structure and diversity advances previous datasets, (ii) we use AndroidControl to quantify how fine-tuning with demonstrations scales when applied to both low- and high-level tasks and to tasks in and out of the domain of the training data, and (iii) we compare fine-tuning to various zero-shot and few-shot baselines, finding that fine-tuning scales favorably in domain, but out of domain, it requires one or two orders of magnitude more data to obtain robust performance on high-level tasks, suggesting that additional approaches may be beneficial for obtaining agents which robustly perform out-of-domain high-level tasks.
@@ 2 Related work
Table 1: Comparison of AndroidControl to existing UI control datasets. We consider device platform, size (as number of task demonstrations), diversity (as number of unique task instructions and apps/websites), average number of steps in a task, how the UI state is captured (UI tree vs. screenshot), and whether tasks are described through high-level instructions or sequences of low-level commands.
@@
Table 1 compares AndroidControl to existing UI control datasets. The structure of these datasets is similar. They consist of a natural language task description and a human-recorded demonstration, in the form of a sequence of UI actions (click, type, swipe, etc.) and associated UI states. What differentiates these datasets is mainly whether they are single-step (as in grounding referring expressions datasets like UIBert [2]), whether the task description is expressed as a high-level goal or as a sequence of low-level step instructions, and how the UI state is represented (screenshot vs. UI tree). Three features make AndroidControl unique. First, for every task, it contains both low-level and high-level instructions generated by human annotators. While a few other datasets contain both these types of annotation, their low-level instructions are either synthetically generated (as in MoTIF [6]) or are limited to one action type (only click actions [22, 33]). In addition to bringing richer language supervision during training, the availability of human-generated low-level instructions allows us to test UI control agents on different levels of task complexity. Second, if we consider the number of unique task instructions and the number of human demonstrations, AndroidControl is the second-largest UI control dataset to date, second only to AitW [27]. [^2] AitW is heavily web focused. If we consider the AitW episodes involving Android apps, we obtain  505k episodes across 167 apps, for a total of 1.6k unique task instructions, a much smaller number of AndroidControl. AitW also does not contain application UI trees, thus making the UI state representations incomplete. The diversity of task scenarios present in AndroidControl is its third differentiating feature: AndroidControl includes tasks from 833 different Android apps, 6 times more than popular datasets like Mind2Web [7] and almost 5 times more than AitW. This diversity makes AndroidControl optimal for realistic, out-of-domain analysis. Note that Mind2Web also provides out-of-domain splits but given its smaller size (2,350 tasks over 137 websites, with a train split of 1k demonstrations) is not suitable for a scaling analysis.
@@
In addition to the datasets listed in Table 1, recent work proposes interactive testing environments for UI control agents [31, 46, 18, 38, 28, 4] where the environment provides the agents with reward signals. These environments are designed for online testing and are limited to no more than 20 applications or websites. The only exception is MiniWob [31] for which task demonstrations have been collected, but the environment consists of much simplified, synthetic websites.
UI control agents
@@
Early UI control agents were trained from scratch using behavioural cloning [15, 21, 22] or reinforcement learning [23, 11]. Current UI agents use pre-trained LLMs and multimodal models. One line of work prompts LLMs in a zero-shot or few-shot regime [39, 27, 12, 18, 17, 44]. Another line of work relies on fine-tuning which is applied end to end [25] or to build specific model capabilities, such as identifying the interactable UI elements in a webpage [43, 10, 7]. To name a few, SeeAct [43], which we use in our evaluation, is a web agent that leverages large multimodal models to understand text and visual elements on webpages. The best-performing SeeAct agent relies on a fine-tuned cross-encoder model to select candidates web elements for interaction. WebGPT [25] fine-tunes GPT-3 to learn to use a web browser. WebAgent [10] pre-trains a T5 model to extract HTML snippets and leverages Flan-U-PaLM to generate Python code to control a web environment. Synapse [44] introduces a trajectory-as-exemplar prompting method where memory of previous interactions allows the agent to perform complex, multi-step tasks.
@@
As evidenced by various LLM studies [5, 16, 13], scaling model and data size for the training leads to steady improvements in domain generalization. On the other hand, when transferring a pre-trained model to a downstream task through fine-tuning, while in-distribution performance improves, a reduction in the robustness to distribution shifts is observed [20, 37, 1]. In this work, we empirically study how scaling data size in fine-tuning affects in-domain and out-of-domain performance of UI control agents. While prior work has tested UI agents using out-of-domain test splits [7, 6], to the best of our knowledge a data scale analysis has not been conducted. To minimize training cost and to maintain the out-of-domain generalization, we evaluate also the option of not fine-tuning, by evaluating multiple zero-shot and few-shot baselines.
@@ 3.1 Data collection
During collection of a demonstration, annotators first provide a high-level description of a task in natural language (e.g., "Add an alarm to wake me up on Saturday mornings at 6am"). We ask annotators to make the descriptions detailed enough to be interpretable without any ambiguity. We also instruct them to always include the name of the target app in the task description, unless obvious (e.g., Google first-party apps such as Clock or Settings). By doing so, the collected data can enable us to test memory-less, single-turn agent interactions.
@@
As both high-level and low-level instructions are generated by annotators, they may contain typo and grammatically errors (for example, the high-level instruction in Figure 1). We left such errors in the dataset because humans often produce errors of this type.
@@ 3.3 Dataset splits
We create a train, a validation and 4 test splits whose number of task demonstrations (episodes) and characteristics are detailed in Table 3. In order to measure how performance scales in domain and out of the domain of the collected data, we create the following test sub-splits: 1) in domain data (IDD): randomly pulled episodes from the same distribution as the training data; 2) app-unseen: a test split using apps not present in the train split; 3) task-unseen: a test split with tasks not present in the train split; and 4) category-unseen: a test split with apps from categories not present in the train split. Note that the test splits may contain overlapping episodes. For example, episodes in the unseen-category split will also be in the unseen-app and unseen-tasks splits. It is also the case that the average number of elements per screen can vary significantly between training and testing. Appendix B.4 explains why that is the case.
@@ 4.1 Agent implementation
We implement a UI control agent for Android. The agent receives task instructions expressed in natural language. It observes the environment (the device) by deriving textual representations of the screen directly from the Android accessibility tree. The screen representation lists the on-screen UI elements. Each element is described according to the following attributes: type, text, content description, bounding boxes and various state tags (e.g., clickable, scrollable, focused, checked). As mobile UI screens may contain hundreds of UI elements (200 on average in AndroidControl, Table 3), we pre-process the screen representations to include only UI elements that have a non-empty text description or UI elements of critical types (switch and edit). This process facilitates the agent’s task and reduces the input’s size. Note that our agent implementation does not directly leverage the page screenshot. While recent work explores how to infer screen representations from raw screens [18], best performance is still reported when using accessibility trees or HTML [38, 43, 28]. We expect the general trends we observe will hold true for multimodal language models.
@@
The agent predicts an action among a set of candidate actions. The set of available actions matches the actions defined by AndroidControl (Table 2) with two main changes. We add a terminate action that the agent predicts when it deems the task complete or infeasible. As this action is not originally provided in the dataset, for training purposes, we artificially insert it at the end of every episode (see Appendix D.1). For efficiency reasons, as in prior work [21], the input_text action is modified to include also the preceding click action necessary to set focus on the target element. The agent predicts the action type and any required arguments for the action, specifically the target UI element for a click action, the text to be typed and the target element for a typing action, the name of an app to open, the direction of a scroll, etc. For an example of screen representation, a summary of the agent’s action space, and more details on the agent implementation please refer to Appendix C.
@@ 4.2 Experimental setup
The LLMs we experiment with include PaLM-2L [9], PaLM-2S [9], Gemini 1.5 Pro [8], GPT-4 and GPT-4 Turbo [26]. We set the temperature to zero for all models to obtain more deterministic responses. To limit compute, we perform fine-tuning only with PaLM-2S, and adopt the parameter efficient tuning approach of LoRA [14]. We set the LoRA rank to 4 when fine-tuning with small amounts of data ($<1$k episodes), while switch to a rank of 64 when using more episodes. For few-shot experiments we use Gemini 1.5 Pro which provides a context window of 1M tokens.
We create SeqIO [29] tasks to extract data points from AndroidControl and to generate prompts and target outputs (more details in Appendix D.1). We setup two SeqIO tasks: (i) SeqIO HL (high-level) where only a high-level instruction is included in the prompt, and (ii) SeqIO LL (low-level) where both a low-level instruction and its corresponding high-level instruction are included. This second task emulates the use case where an LLM is used for decomposing high-level instructions into a sequence of low-level commands and another LLM is used for grounding; the assumption is that this grounding LLM may improve performance by having access to the context of the high-level command. In addition to the natural language instruction(s), each data point contains the textual description of the start screen, the history of performed actions, and the ground-truth action. Through these two SeqIO tasks, we investigate how a model performs on simpler (LL) or harder (HL) task instructions.
@@
We test four zero-shot methods. (i) We use the AitW [27] prompt, specifically designed for Android and the PaLM model, without any modifications. (ii) We adapt the best-performing SeeAct [43] variant ("choice") which grounds actions via textual choices. SeeAct was originally designed for GPT-4V for web navigation tasks. At each step, SeeAct queries the LLM twice. In the first query, it analyzes the state and performs reasoning. Then, in the second query, it asks the LLM to select an action from multiple choices. We use the SeeAct prompt by Rawles et al. [28] adapted to work on mobile and to take textual representations of Android screens as input. (iii) We evaluate the text-only version of M3A [28] that combines ReAct-style [42] and Reflexion-style [32] prompting. (iv) Finally, we test a zero-shot prompt (implementation in Appendix D.2.2) of the same form we use with our agent described above. This allows us to measure performance of a base model for our agent without any fine-tuning. This prompt emphasizes the use of a screen description composed of UI elements, hence the name ER (Element Representations). Note that with the exception of ER, which we ran with all 4 base models, to limit prompt sensitivity [30], we ran the other prompts with the model family they were originally designed for.
@@
Table 4: Performance on the IDD sub-split of Random-500. For few-shot (FS) and LoRA-tuned (LT) methods -$X$ (-$all$) indicates $X$ ($all$) episodes are used in the prompt or in training. Unless noted with a “r64” for models fine-tuned with a LoRA rank of 64, LoRA-tuned models use rank=4.